# Investigating the heterogenous effects of a massive content moderation intervention via Difference-in-Differences

### v3

@@ 4.2 RQ2: Extreme user reactions to The Great Ban
Figure 3: Subreddit-wise percentage toxicity increment (orange bars, right-hand side) and total number of users (purple bars, left-hand side), for the banned and baseline (non-banned) subreddits. Subreddits where the toxicity increases outweight the decreases are highlighted in red, while the others in cyan.
Figure 4: Subreddit-wise percentage toxicity increment of the outliers users (orange bars, right-hand side) and percentage of outlier users out of all subreddit users (purple bars, left-hand side), for the banned and baseline (non-banned) subreddits. Subreddits where the toxicity increases of the outlier users outweight the decreases are highlighted in red, while the others in cyan. The subreddit r/hatecrimehoaxes is black-colored since it does not contain any outlier.
@@
Figure 3 illustrates the percentage increment toxicity for the banned and baseline subreddits, along with the number of users in those subreddits. The corresponding decrease in toxicity, which is relative to the left tail, can be easily computed as $\Delta t^{-}=100-\Delta t^{+}$. By looking at the percentage toxicity increment for the banned subreddits, we can deduce that the contributions of the left and right tails are quite balanced across nearly all subreddits. For example, in the case of r/the_donald these contributions are perfectly balanced. This is also the same for the baseline subreddits. For 13 out of 15 banned subreddits (blue colored), the decrease in toxicity slightly exceeds the increase, which leads to the overall modest reduction in toxicity observed in RQ1. However, we can observe the opposite trend in r/debatealtright and r/imgoingtohellforthis2 (red colored).
@@
To further assess the behavior of outlier users – those who drastically increased or decreased their toxicity – we recompute $\Delta t^{+}$ and $\Delta t^{-}$ by only considering those users whose $|\Delta t(i)|>0.25$. Figure 4 reiterates the previous analysis for outlier users. The proportion of outlier users (left-hand side of the figure) remains relatively small in all subreddits, with r/gendercritical exhibiting the highest percentage with almost 6% of users, while r/hatecrimehoaxes does not have outliers at all. The results are notably different than those of Figure 3, as for 11 out of 15 banned subreddits (red colored) the increase in toxicity (right-hand side of the figure) significantly outweighs the decrease. This does not apply to r/gendercritical, r/soyboys, and r/oandaexlusiveforms, as for the first two subreddits, the increase in toxicity is less than 50% and no increase is observed for the latter. For what concerns the baseline subreddits, the vast majority of outliers exhibited a marked toxicity increment. However, the number of outliers within the baseline subreddits is much lower with respect to the outliers in the banned subreddits, and ranges from 0.13% to 0.22%. In addition to these results, we also repeated the analysis for multiple values of $|\Delta t(i)|$, as described in Cima et al. (2024b). We found that the obverved behavior is independent of the way in which outlier users are defined, which supports the robusteness of our finding. Overall, our results reveal that the phenomenon of outlier users who drastically increased their toxicity post-ban is peculiar of the banned subreddits, and can possibly be explained as a negative reaction of such users to the ban itself. Moreover, our findings also underscore that extremely toxic users are spread across almost all banned subreddits, which is a sign of a systemic reaction to the ban.
4.3 RQ3: Behavioral dynamics of extremely toxic users
@@
Subreddits participation. Initially, we explore the subreddit participation habits of these users as a function of their change in toxicity. This analysis aims to ascertain whether each of the most toxic users engaged in a few or many of the banned subreddits. Figure 5 suggests that only a small number of users participated in multiple subreddits. Instead, the vast majority of users participated only in a few subreddits, typically ranging between one and three, consistently with the echo chamber theory (Cinelli et al., 2021; Valensise et al., 2023). Moreover, user toxicity changes do not depend on the number of subreddits in which the user participated.
@@
Activity. Next, we look into the relationship between changes in toxicity and activity. We measure user activity as the number of posted comments. Conjointly studying user activity and toxicity is important since toxic users pose a meaningful threat only when high toxicity coincides with elevated activity, which holds the potential to cause a substantial impact on the platform. This joint analysis also enables the identification of those users who are both highly toxic and remarkably active, allowing to prioritize moderation interventions towards them (Cresci et al., 2022). The left-most plot of Figure 6 presents a density scatterplot of the changes in toxicity and activity for each considered user. The plot area is divided into four quadrants, corresponding to the combinations of increase/decrease in toxicity/activity. As shown, the bulk of the distribution lies close to the plot origin, meaning that the ban caused minor changes in both toxicity and activity. Moreover, those users who experienced marked changes in toxicity only had negligible changes in activity, and vice versa. The insets in Figure 6 show the number of users laying in each quadrant of the scatterplots. The most frequent reaction to the ban was a reduction in both toxicity and activity (38% users lay in Q3). This is also evident in the left-most plot of Figure 7, which pictures the relationship between mean toxicity and user activity, after the ban. It shows how the vast majority of users who exhibit low levels of toxicity are moderately active. The opposite behavior is instead the least frequent (16% users are in Q1). Overall this analysis suggests that the resentful users might have had a limited impact on the platform due to their low activity.
@@
Engagement. Finally, we examine the relationship between changes in toxicity and obtained engagement. We measure the engagement obtained by each comment as its Reddit score, computed as the difference between the upvotes and downvotes to that comment. Positive scores denote a positive engagement of the community with a comment, while negative scores generally denote disagreement. The engagement of a user is the mean engagement obtained by its comments. This joint analysis of toxicity and engagement is relevant as it allows an understanding of whether post-ban toxic comments are tolerated, and possibly even encouraged, on Reddit. The right-most plot of Figure 6 shows the density scatterplot of this relationship while the right-most plot of Figure 7 shows the distribution of mean toxicity with respect to the mean engagement score after the ban. Similarly to what occurred for activity in Figure 6 and Figure 7, we found that the vast majority of users experienced modest changes in toxicity and engagement, and lower values of engagement scores correspond to lower levels of toxicity. Again, the most frequent reaction was a simultaneous reduction of both toxicity and obtained engagement (31% users are in Q3) and the least frequent one was a simultaneous increase of both (21% users are in Q1).
Lastly, we analyze the relationship between toxicity and engagement also before the ban, by investigating differences between the engagement obtained within vs outside of the 15 banned subreddits. Figure 8 provides the results of this comparison, highlighting that toxic comments received significantly more engagement within the banned subreddits than outside of them (i.e., in the rest of Reddit). This result corroborates that of the right-most plot of Figure 6 in that it reinforces the idea that toxicity is not encouraged on Reddit, with the exception of a few toxic subreddits.
Figure 8: The difference in scores between toxic comments made within and outside the 15 banned subreddits as a function of comment toxicity. The significance of these differences is determined using the t-test for unpaired data: *: $p<0.1$; **: $p<0.05$; ***: $p<0.01$.
@@ 5 Discussion
Our study provides insights into the consequences of The Great Ban, a crucial example of deplatforming that affected over 2,000 subreddits. We highlight that 15.6% of the users involved abandoned the platform, however those who remained reduced their average toxicity by 4.1%. At the same time, however, approximately 5% of all users significantly escalated their toxicity. The presence of these so-called resentful users was observed across many of the analyzed subreddits. However, only 16% of users increased both their toxicity and activity levels, and just 21% of users who became more toxic also received positive feedback from other users. Our findings provide new insights regarding the impact of The Great Ban and, in general, on the heterogeneous and unintended negative reactions to moderation interventions (Trujillo and Cresci, 2023). Additionally, they shed light on several aspects regarding the design and implementation of effective moderation strategies, underlining the inherent challenges in moderating online platforms.
@@
Effectiveness of the moderation. In the current literature, the effectiveness of content moderation has typically been evaluated by examining the changes in these interventions in terms of activity and toxicity  (Chandrasekharan et al., 2017; Jhaver et al., 2021). In this context, our study showed that a significant portion of toxic users left the platform, while those who remained exhibited a modest decrease in toxicity along with a notable decline in activity. Apparently, these results suggest successful moderation. However, it is crucial to carefully examine these findings by considering potential unintended consequences. For instance, the deplatforming of toxic users may move the issue somewhere else rather than solve it. As highlighted by recent studies, moderated users might migrate to other platforms (Horta Ribeiro et al., 2021), possibly indicating that their toxic interactions may have been displaced rather than mitigated. Moreover, migrated users may engage in much more toxic and aggressive behavior in the new platforms (Horta Ribeiro et al., 2021). Finally, user abandonment and reduced activity after the ban could represent a significant issue for Reddit, as user engagement and interactions are the main sources of revenue for online platforms (Trujillo and Cresci, 2022; Tessa et al., 2024). As a result, the presumed effectiveness of The Great Ban in reducing toxicity should be interpreted carefully as it can negatively impact not only the platform’s revenue because of abandonment or reduced engagement, but also outside Reddit because of user migration. Future efforts should focus on finding a balance between developing strategies that effectively reduce toxicity while minimizing user abandonment or decline in activity (Tessa et al., 2024). In fact, effective moderation should foster healthier online communities while safeguarding platform safety and economic sustainability. Otherwise, platforms may be discouraged from implementing rigorous moderation measures.
@@
Divergent reactions to moderation. Our study showed that, despite a slight decrease in toxicity, a significant minority of users exhibited substantial increases in their toxic behavior. This finding has significant implications for evaluating and designing moderation interventions. First, it underscores the complexity of user responses to content moderation, an area that is still vastly underexplored and warrants further investigation (Jhaver et al., 2023). Additionally, it brings to light the necessity for personalization in the context of content moderation. A broad intervention like The Great Ban – which impacted thousands of subreddits and tens of thousands of users – may not adequately fit the different motivations and personalities of those affected. Here, this was shown by the minority of resentful users who significantly escalated their toxic behavior. Understanding the aspects that lead to these divergent responses is crucial for developing effective moderation interventions. Future research and on-field applications should focus on user profiling, by taking into account individual personal traits, past behavior, and contextual factors to effectively customize moderation interventions (Cresci et al., 2022; Annamoradnejad et al., 2022). In addition, user migration combined to a high percentage of users displaying increased toxicity raises concerns about the potential radicalization effects of moderation and its inadvertent role in amplifying echo chambers (Horta Ribeiro et al., 2021). This indicates that some users may not respond positively to specific moderation interventions, potentially resulting in more radical behaviors. Our findings emphasize the careful balance needed in content moderation, which not only should locally reduce toxicity but also should prevent unintended consequences that may lead to polarization or radicalization within specific user groups.
@@
Impact of resentful users. The finding that the vast majority of resentful users – who exhibited increased toxicity following the ban – did not concurrently escalate their activity offers a novel perspective on the little-studied interplay of toxicity and user activity. Contrary to the suspicion that highly toxic users might also be highly active, our results underscore a limited overlap between these two dimensions of user behavior. This suggests that a considerable portion of users expressing heightened toxicity post-ban did not translate their resentment into increased participation on the platform. In turn, this result partly mitigates the previous concerns about the divergent and adverse reactions to The Great Ban. Furthermore, the observation that users who intensified their toxic behaviors did not receive heightened positive social feedback from their peers resonates with some recent research on the drivers of online hateful and toxic speech. The latter posits that online toxicity may be driven by the pursuit of social approval rather than by a desire to harm (Jiang et al., 2023). To this end, our results demonstrate that the Reddit community did not endorse or support the increased toxicity exhibited by these users. According to the aforementioned theories, toxic users who do not receive the expected levels of social approval should decrease their toxicity over time (Jiang et al., 2023). However, we measured overall stationary toxicity trends over the course of the 7 months after the ban. Our results thus apparently challenge the social approval theory, in support of competing theories according to which at least a subset of toxic and hateful users seek amusement by harassing others, as in trolling behavior (Chandrasekharan et al., 2022; Uyheng et al., 2022). On the one hand, these results call for additional research to identify the types and characteristics of toxic users. On the other hand, they reinforce the need for nuanced and diversified moderation strategies that are capable of addressing the complexity and multiplicity of online user behavior.
Limitations and Future works. In this study, we make use of a large dataset consisting of comments posted by 34K Reddit users during 7 months before and after The Great Ban took place. This means that our findings may be limited to Reddit and the effects of The Great Ban itself. Moreover, online platforms are dynamic and characterized by frequent changes in user behavior, community norms, and platform policies. Our study takes into account a specific timeframe and while it covers a relatively extended period, it limits the generalizability of our findings to different periods of time and somewhat constrains our ability to fully account for long-term or evolving trends. In addition, our evaluation of the toxicity scores could be biased because of the use of Detoxify, a machine learning model. The observational nature of our work constitutes another limitation, as it limits our ability to consider external events that might have independently influenced user behavior outside of The Great Ban. To conclude, data on user demographics, motivations, and contextual factors is missing from our dataset. Having this type of information could have offered a more refined interpretation of our results. Future research could look more into the unexplored connections between user characteristics and moderation outcomes, serving as an initial step towards the implementation of more targeted and personalized moderation interventions (Cresci et al., 2022; Tessa et al., 2024).

### v4

@@ 4.2 RQ2: Extreme user reactions to The Great Ban
Figure 3: Subreddit-wise percentage toxicity increment (orange bars, right-hand side) and total number of users (purple bars, left-hand side), for the banned and baseline (non-banned) subreddits. Subreddits where the toxicity increases outweight the decreases are highlighted in red ($\uparrow$), while the others in cyan ($\downarrow$).
Figure 4: Subreddit-wise percentage toxicity increment of the outliers users (orange bars, right-hand side) and percentage of outlier users out of all subreddit users (purple bars, left-hand side), for the banned and baseline (non-banned) subreddits. Subreddits where the toxicity increases of the outlier users outweight the decreases are highlighted in red ($\uparrow$), while the others in cyan ($\downarrow$). The subreddit r/hatecrimehoaxes is black-colored since it does not contain any outlier.
@@
Figure 3 illustrates the percentage increment toxicity for the banned and baseline subreddits, along with the number of users in those subreddits. The corresponding decrease in toxicity, which is relative to the left tail, can be easily computed as $\Delta t^{-}=100-\Delta t^{+}$. By looking at the percentage toxicity increment for the banned subreddits, we can deduce that the contributions of the left and right tails are quite balanced across nearly all subreddits. For example, in the case of r/the_donald these contributions are perfectly balanced. This is also the same for the baseline subreddits. For 13 out of 15 banned subreddits (cyan colored), the decrease in toxicity slightly exceeds the increase ($\downarrow$), which leads to the overall modest reduction in toxicity observed in RQ1. However, we can observe the opposite trend in r/debatealtright and r/imgoingtohellforthis2 (red colored $\uparrow$). By repeating the same analysis by using the frequency of toxic comments instead of the mean toxicity, we obtained results closely matching those in Figure 3, showing robustness to the choice of a toxicity indicator.
@@
To further assess the behavior of outlier users – those who drastically increased or decreased their toxicity – we recompute $\Delta t^{+}$ and $\Delta t^{-}$ by only considering those users whose $|\Delta t(i)|>0.25$. Figure 4 reiterates the previous analysis for outlier users. The proportion of outlier users (left-hand side of the figure) remains relatively small in all subreddits, with r/gendercritical exhibiting the highest percentage with almost 6% of users, while r/hatecrimehoaxes does not have outliers at all. The results are notably different than those of Figure 3, as for 11 out of 15 banned subreddits (red colored $\uparrow$) the increase in toxicity (right-hand side of the figure) significantly outweighs the decrease. This does not apply to r/gendercritical, r/soyboys, and r/oandaexlusiveforms, as for the first two subreddits, the increase in toxicity is less than 50% and no increase is observed for the latter. For what concerns the baseline subreddits, the vast majority of outliers exhibited a marked toxicity increment. However, the number of outliers within the baseline subreddits is much lower with respect to the outliers in the banned subreddits, and ranges from 0.13% to 0.22%. In addition to these results, we also repeated the analysis for multiple values of $|\Delta t(i)|$, as described in [13]. We found that the obverved behavior is independent of the way in which outlier users are defined, which supports the robusteness of our finding. Overall, our results reveal that the phenomenon of outlier users who drastically increased their toxicity post-ban is peculiar of the banned subreddits, and can possibly be explained as a negative reaction of such users to the ban itself. Moreover, our findings also underscore that extremely toxic users are spread across almost all banned subreddits, which is a sign of a systemic reaction to the ban.
4.3 RQ3: Relationship between toxicity, activity, and engagement
@@
Subreddits participation. Initially, we explore the subreddit participation habits of these users as a function of their change in toxicity. This analysis aims to ascertain whether each of the most toxic users engaged in a few or many of the banned subreddits. Figure 5 suggests that only a small number of users participated in multiple subreddits. Instead, the vast majority of users participated only in a few subreddits, typically ranging between one and three, consistently with the echo chamber theory [14, 61]. Moreover, user toxicity changes do not depend on the number of subreddits in which the user participated.
@@
Activity. Next, we look into the relationship between changes in toxicity and activity. We measure user activity as the number of posted comments. Conjointly studying user activity and toxicity is important since toxic users pose a meaningful threat only when high toxicity coincides with elevated activity, which holds the potential to cause a substantial impact on the platform. This joint analysis also enables the identification of those users who are both highly toxic and remarkably active, allowing to prioritize moderation interventions towards them [17]. The left-most plot of Figure 6 presents a density scatterplot of the changes in toxicity and activity for each considered user. The plot area is divided into four quadrants, corresponding to the combinations of increase/decrease in toxicity/activity. As shown, the bulk of the distribution lies close to the plot origin, meaning that the ban caused minor changes in both toxicity and activity. Moreover, those users who experienced marked changes in toxicity only had negligible changes in activity, and vice versa. The insets in Figure 6 show the number of users laying in each quadrant of the scatterplots. The most frequent reaction to the ban was a reduction in both toxicity and activity (38% users lay in Q3). This is also evident in the left-most plot of Figure 7, which pictures the relationship between mean toxicity and user activity, after the ban. It shows how the vast majority of users who exhibit low levels of toxicity are moderately active. The opposite behavior is instead the least frequent (16% users are in Q1). Overall this analysis suggests that the resentful users might have had a limited impact on the platform due to their low activity.
@@
Engagement. Finally, we examine the relationship between changes in toxicity and obtained engagement. We measure the engagement obtained by each comment as its Reddit score, computed as the difference between the upvotes and downvotes to that comment. Positive scores denote a positive engagement of the community with a comment, while negative scores generally denote disagreement. The engagement of a user is the mean engagement obtained by its comments. This joint analysis of toxicity and engagement is relevant as it allows an understanding of whether post-ban toxic comments are tolerated, and possibly even encouraged, on Reddit. The right-most plot of Figure 6 shows the density scatterplot of this relationship while the right-most plot of Figure 7 shows the distribution of mean toxicity with respect to the mean engagement score after the ban. Similarly to what occurred for activity in Figures 6 and 7, we found that the vast majority of users experienced modest changes in toxicity and engagement, and lower values of engagement scores correspond to lower levels of toxicity. Again, the most frequent reaction was a simultaneous reduction of both toxicity and obtained engagement (31% users are in Q3) and the least frequent one was a simultaneous increase of both (21% users are in Q1).
Lastly, we analyze the relationship between toxicity and engagement before the ban, by investigating differences between the engagement obtained within vs outside of the 15 banned subreddits. Figure 8 provides the results of this comparison, which shows that toxic comments received slightly more engagement within the banned subreddits than in the rest of Reddit. This is evident from the higher concentration of toxic content in these communities, where the toxicity score distribution is more skewed toward large values. Additionally, comments in the banned subreddits received significantly higher mean scores than those elsewhere ($p<0.01$), suggesting that engagement with toxic content was strong and that such content was not actively discouraged. We further examined the correlation between toxicity and user-assigned scores, finding very weak yet opposite trends: a positive Pearson correlation ($r=+0.007$) in banned subreddits and a negative one ($r=-0.006$) in the rest of Reddit. While not statistically significant, this pattern suggests that toxicity might have been tolerated or even slightly incentivized within banned communities, whereas it was mildly discouraged elsewhere. This result corroborates that of the right-most plot of Figure 6 in that it reinforces the idea that toxicity is not encouraged on Reddit, except for a few toxic subreddits.
Figure 8: Density scatterplots of the relationships between toxicity and score within the banned subreddits (left-hand side) and in the rest of Reddit (right-hand side). For each scatterplot, we also report the Pearson correlation $r$ between toxicity and score, and the mean score.
@@ 5 Discussion
Our study provides insights into the consequences of The Great Ban, a crucial example of deplatforming that affected over 2,000 subreddits. We highlight that 15.6% of the users involved abandoned the platform, however those who remained reduced their average toxicity by 4.1%. At the same time, however, approximately 5% of all users significantly escalated their toxicity. The presence of these so-called resentful users was observed across many of the analyzed subreddits. However, only 16% of users increased both their toxicity and activity levels, and just 21% of users who became more toxic also received positive feedback from other users. Our findings provide new insights regarding the impact of The Great Ban and, in general, on the heterogeneous and unintended negative reactions to moderation interventions [57]. Additionally, they shed light on several aspects regarding the design and implementation of effective moderation strategies, underlining the inherent challenges in regulating online platforms.
5.1 Effectiveness of the moderation
@@
In the current literature, the effectiveness of content moderation has typically been evaluated by examining the changes occurred after certain moderation interventions in terms of activity and toxicity [10, 34]. In this context, our study showed that a significant portion of toxic users left the platform, while those who remained exhibited a modest decrease in toxicity along with a notable decline in activity. Apparently, these results suggest successful moderation. However, it is crucial to carefully examine these findings by considering potential unintended consequences. For instance, the deplatforming of toxic users may move the issue somewhere else rather than solve it. As highlighted by recent studies, moderated users might migrate to other platforms [31], possibly indicating that their toxic interactions may have been displaced rather than mitigated. Moreover, migrated users may engage in much more toxic and aggressive behavior in the new platforms [31]. Finally, user abandonment and reduced activity after the ban could represent a significant issue for Reddit, as user engagement and interactions are the main sources of revenue for online platforms [56, 55]. As a result, the presumed effectiveness of The Great Ban in reducing toxicity should be interpreted carefully as it can negatively impact not only the platform’s revenue because of abandonment or reduced engagement, but also outside Reddit because of user migration. Future efforts should focus on finding a balance between developing strategies that effectively reduce toxicity while minimizing user abandonment or decline in activity [55]. In fact, effective moderation should foster healthier online communities while safeguarding platform safety and economic sustainability. Otherwise, platforms may be discouraged from implementing rigorous moderation measures.
5.2 Divergent reactions to moderation
@@
Our study showed that, despite a slight decrease in toxicity at the aggregate level, a significant minority of users exhibited substantial increases in their toxic behavior. This finding has important implications for evaluating and designing moderation interventions. First, it underscores the complexity of user responses to content moderation, an area that remains vastly underexplored and warrants further investigation [35]. Additionally, it brings to light the necessity for personalization in the context of content moderation [17]. A broad intervention like The Great Ban – which affected thousands of subreddits and tens of thousands of users – may not adequately address the diverse motivations and behavioral patterns of those impacted. Here, this was demonstrated by the subset of users who reacted with markedly increased toxicity. Our analyses confirm that these escalations were not artifacts of sparse post-ban activity or indicator instability: users who became more toxic often did so both in terms of average toxicity and in the absolute number of toxic comments. This reinforces the interpretation that a genuine shift occurred in the behavior of a small segment of users. Understanding what drives such divergent reactions is crucial for developing more nuanced and effective moderation interventions. Future research and real-world implementations should therefore focus on user profiling, taking into account individual traits, behavioral history, and contextual factors to support more adaptive and personalized moderation strategies [12, 2]. In addition, the observed user migration – combined with a sizable fraction of users displaying increased toxicity – raises concerns about the potential radicalization effects of moderation and its unintended role in reinforcing echo chambers [31]. These findings suggest that some users may not respond positively to specific forms of intervention, and may instead engage in more extreme behaviors. Overall, our results emphasize the careful balance needed in content moderation: interventions should aim not only to reduce toxicity at the community level but also to anticipate and mitigate the risk of backlash among vulnerable user groups.
5.3 Impact of resentful users
@@
The finding that the vast majority of resentful users – who exhibited increased toxicity following the ban – did not concurrently escalate their activity offers a novel perspective on the little-studied interplay of toxicity and user activity. Contrary to the suspicion that highly toxic users might also be highly active, our results underscore a limited overlap between these two dimensions of user behavior. This suggests that a considerable portion of users expressing heightened toxicity post-ban did not translate their resentment into increased participation on the platform. In turn, this result partly mitigates the previous concerns about the divergent and adverse reactions to The Great Ban. Furthermore, the observation that users who intensified their toxic behaviors did not receive heightened positive social feedback from their peers resonates with some recent research on the drivers of online hateful and toxic speech. The latter posits that online toxicity may be driven by the pursuit of social approval rather than by a desire to harm [36]. To this end, our results demonstrate that the Reddit community did not endorse or support the increased toxicity exhibited by these users. According to the aforementioned theories, toxic users who do not receive the expected levels of social approval should decrease their toxicity over time [36]. However, we measured overall stationary toxicity trends over the course of the 7 months after the ban. Our results thus apparently challenge the social approval theory, in support of competing theories according to which at least a subset of toxic and hateful users seek amusement by harassing others, as in trolling behavior [9, 60]. On the one hand, these results call for additional research to identify the types and characteristics of toxic users. On the other hand, they reinforce the need for nuanced and diversified moderation strategies that are capable of addressing the complexity and multiplicity of online user behavior [17].
5.4 Legal and regulatory implications
Our findings reveal the moderate effectiveness of The Great Ban. While the intervention led to a mild overall reduction in toxicity, it also resulted in heterogeneous effects: a subset of users became significantly more toxic, while a considerable portion of previously active users abandoned the platform altogether. These mixed outcomes align with broader findings in the literature, suggesting that moderation interventions do not always produce the intended or necessary effects [5, 43, 31, 56]. These results raise critical questions about the effectiveness and accountability of content moderation policies, particularly in the context of evolving regulatory frameworks in both the US and Europe.
In the US, our results connect to ongoing debates surrounding Section 230 of the Communications Decency Act [1], which provides platforms with broad discretion in how they moderate content. Notably, Section 230 does not require platforms to moderate content at all, nor does it hold them accountable for the effectiveness of their moderation efforts. The fact that The Great Ban produced only marginal improvements in reducing toxicity, while simultaneously exacerbating toxicity in a subset of users, underscores a key issue in the current legal landscape: platforms are not obligated to implement moderation strategies that are evidence-based or demonstrably effective. These findings contribute to recent discussions on potential reforms to Section 230, which aim to increase platform responsibility for harmful content [3, 15]. Such reforms could incentivize the development of more strategic, proactive, and adaptive moderation frameworks [27, 55], shifting away from reactive mass-deplatforming toward targeted interventions that minimize unintended consequences [17, 12].
On the other hand, the European Union’s Digital Services Act (DSA) takes a markedly different approach by imposing stricter transparency and accountability requirements on content moderation [23]. Our findings – particularly the observation that a fraction of users became more toxic and resentful following the ban – relate to key DSA provisions that mandate platforms to assess the risks of their moderation policies, including unintended side effects, before enforcement. While this requirement represents an important step toward more responsible platform governance, significant uncertainties remain regarding: (i) how platforms currently evaluate the risks associated with their moderation policies, (ii) whether they possess the capacity to effectively measure or mitigate these risks, and (iii) the extent to which opaque moderation decisions impact users and broader platform dynamics. These challenges call for renewed multidisciplinary research efforts to better assess the unintended consequences of content moderation and to develop evidence-based mitigation strategies that align with evolving regulatory expectations [55, 12].
In conclusion, our study highlights how the current regulatory landscape of content moderation varies significantly across regions. The US model prioritizes platform autonomy, allowing platforms to moderate at their own discretion without clear accountability for their outcomes. In contrast, EU regulations emphasize transparency, risk assessment, and harm mitigation, pushing platforms to justify and evaluate their moderation decisions more rigorously [58, 38, 54, 51]. As global challenges in platform governance continue to evolve, platforms may increasingly need to tailor their moderation strategies to comply with diverging regional requirements – a shift that could shape the future of online content regulation worldwide.
5.5 Limitations and Future Works
In this study, we make use of a large dataset consisting of comments posted by 34K Reddit users during seven months before and after The Great Ban. This means that our findings may be limited to Reddit and the specific circumstances of this intervention. Moreover, online platforms are dynamic and characterized by frequent changes in user behavior, community norms, and platform policies. Our study takes into account a fixed timeframe – while relatively extended, it constrains the generalizability of our findings to other periods and evolving trends. In addition, our evaluation of toxicity relies on Detoxify, a machine-learning classifier whose output can be sensitive to the volume of comments per user. To guard against this potential artifact, we conducted robustness checks using alternative toxicity measures (e.g., proportion and absolute count of toxic comments) and restricted analyses to users with stable, substantial activity. These additional experiments confirmed that our main conclusions – namely, a modest overall toxicity decline alongside a small subset of users who escalated their toxic behavior – hold under different definitions and activity constraints. The observational nature of our work also constitutes a limitation, as it reduces our ability to account for external events that might have independently influenced user behavior. For instance, major socio-political events such as the George Floyd protests in mid-2020, which sparked widespread discussions on racism and online hate speech [44], may have contributed to shifts in both user behavior and moderation enforcement. Likewise, the COVID-19 pandemic, which heavily influenced online interactions and engagement patterns during 2020, could represent another confounding factor. While our Difference-in-Differences (DiD) design and inclusion of a baseline group mitigate such concerns to some extent, we acknowledge that these external factors may have shaped the observed behavioral changes. Finally, data on user demographics, motivations, and contextual factors is missing from our dataset. Such information could have offered a more granular interpretation of our results. Future work should explore the connections between user characteristics and moderation outcomes, supporting the development of more targeted, personalized interventions [17, 55, 12]. Moreover, our methodology for selecting the baseline subreddits introduces an additional limitation. Our keyword-based approach tends to favor larger subreddits, which may introduce a size-related bias compared to the typically smaller banned communities. While this choice ensures greater statistical stability and robustness of observed trends, it may reduce thematic comparability. Alternative options—such as smaller, thematically closer subreddits—were considered but often violate the Stable Unit Treatment Value Assumption (SUTVA) due to overlap with the treated groups, thereby compromising causal validity. Future research could further explore this trade-off by identifying alternative reference groups based on user demographics, interaction patterns, or pre-ban engagement similarity, to better balance topical relevance and statistical independence.
5.6 Ethical considerations